# Self-Attention Limits Working Memory Capacity of Transformer-Based Models

### v1

@@ Abstract
Recent work on Transformer-based large language models (LLMs) has revealed striking limits in their working memory capacity, similar to what has been found in human behavioral studies. Specifically, these models’ performance drops significantly on N-back tasks as N increases. However, there is still a lack of mechanistic interpretability as to why this phenomenon would arise. Inspired by the executive attention theory from behavioral sciences, we hypothesize that the self-attention mechanism within Transformer-based models might be responsible for their working memory capacity limits. To test this hypothesis, we train vanilla decoder-only transformers to perform N-back tasks and find that attention scores gradually aggregate to the N-back positions over training, suggesting that the model masters the task by learning a strategy to pay attention to the relationship between the current position and the N-back position. Critically, we find that the total entropy of the attention score matrix increases as N increases, suggesting that the dispersion of attention scores might be the cause of the capacity limit observed in N-back tasks.
@@ 1 Introduction
However, due to the black-box nature of LLMs, we still lack mechanistic insights as to why the observed capacity limit would emerge, especially given the fact that the length of N-back task sequences (e.g., 24 letters in [9]) is well within the context length of these models [16]. To answer this question, we were inspired by the executive attention theory [7, 5, 6] in human working memory research. The executive attention theory proposes that working memory requires executive attention to maintain access to information in the face of interference. suggesting that it is the scarcity of attentional resources [12, 14], but not memory storage itself, that is responsible for working memory capacity limits. In Transformer-based LLMs, the self-attention mechanism computes the importance of each element in the input sequence relative to other elements. While this approach allows the model to focus on relevant information, as N increases in the N-back task, it could be increasingly hard to maintain focus between distant positions. Therefore, we hypothesize that self-attention might be the cause of working memory capacity limits in Transformer-based models.
@@ 2 Methods
We use vanilla Transformers in order to facilitate interpretability, as done in prior work aiming to better understand computations in Transformers in more controlled task settings [4, 13]. We mainly focus our analysis on a causal Transformer containing one decoder layer with only one attention head (Figure 6 in Appendix), although we also test a few architectural variants in the number of decoder layers ($\mathsf{L}$) and number of attention heads per layer ($\mathsf{H}$) for comparisons (see Section 3 for details). The decoder layer contains masked self-attention so that for each position in the sequence the model can only attend to the current and previous positions. No multi-layer feed-forward networks or layer normalization are applied. The decoder layer is then followed by an unembedding layer to project the decoder outputs to two logits (representing match and nonmatch) for each position.
@@ 3 Results
Figure 3: the model learns to attend target locations over training epochs. Here we show attention maps of a 1-layer, 1-head Transformer model trained on the 3-back task as an example. See Appendix for attention maps in the 1-back and 2-back tasks.
@@
Figure 4: (a)-(c): the relationship between test accuracy at position $i$ and the attention score at position $i-N$. Colors represent different epochs each dot belongs to. (d)-(f): same as (a)-(c) but colors represent different position each dot belongs to.
@@
To further investigate the relationship between attention scores and test accuracy, we plot accuracy at position $i$ against the attention score at the position $i-N$ over training epochs ($i\in\{1..24\}$, $N\in\{1,2,3\}$). The accuracy at position $i$ is defined as the percentage of the model making a correct prediction at position $i$. Over training epochs, we find that the attention score at position $i-N$ increases along with the accuracy at position $i$ (Figure 4a-c). We reason that in order to produce an accurate prediction at position $i$, the Transformer model needs to learn to put most attention on the $i-N$ position and reduce dispersion of attention to other positions. To better visualize dispersion of attention scores across positions, we use the same data in Figure 4a-c but assign colors to the dots according to which position each dot belongs to (Figure 4d-f). This reveals a clear pattern that attention scores get dispersed at later locations, suggesting that more interference is caused when there are more preceding positions.
@@
We find that $H_{N}$ increases as N increases, leading to the decrease in test accuracy (Figure 5). We infer that as N increases, it gets harder for the model to learn to attend to the N-back letter and the model is less confident about which letter is important, leading to higher entropy and lower accuracy.
4 Discussion

### v2

@@ Abstract
Recent work on Transformer-based large language models (LLMs) has revealed striking limits in their working memory capacity, similar to what has been found in human behavioral studies. Specifically, these models’ performance drops significantly on N-back tasks as N increases. However, there is still a lack of mechanistic interpretability as to why this phenomenon would arise. Inspired by the executive attention theory from behavioral sciences, we hypothesize that the self-attention mechanism within Transformer-based models might be responsible for their working memory capacity limits. To test this hypothesis, we train vanilla decoder-only transformers to perform N-back tasks and find that attention scores gradually aggregate to the N-back positions over training, suggesting that the model masters the task by learning a strategy to pay attention to the relationship between the current position and the N-back position. Critically, we find that the total entropy of the attention score matrix increases as N increases, suggesting that the dispersion of attention scores might be the cause of the capacity limit observed in N-back tasks. Our findings thus offer insights into the shared role of attention in both human and artificial intelligence. Moreover, the limitations of the self-attention mechanism revealed in the current study could inform future efforts to design more powerful model architectures with enhanced working memory capacity and cognitive capabilities.
@@ 1 Introduction
However, due to the black-box nature of LLMs, we still lack mechanistic insights as to why the observed capacity limit would emerge, especially given the fact that the length of N-back task sequences (e.g., 24 letters in [9]) is well within the context length of these models [16]. To answer this question, we were inspired by the executive attention theory [7, 5, 6] in human working memory research. The executive attention theory proposes that working memory requires executive attention to maintain access to information in the face of interference, suggesting that it is the scarcity of attentional resources [12, 14], but not memory storage itself, that is responsible for working memory capacity limits. In Transformer-based LLMs, the self-attention mechanism computes the importance of each element in the input sequence relative to other elements. While this approach allows the model to focus on relevant information, as N increases in the N-back task, it could be increasingly hard to maintain focus between distant positions. Therefore, we hypothesize that self-attention might be the cause of working memory capacity limits in Transformer-based models.
@@ 2 Methods
We use vanilla Transformers in order to facilitate interpretability, as done in prior work aiming to better understand computations in Transformers in more controlled task settings [4, 13]. We mainly focus our analysis on a causal Transformer containing one decoder layer with only one attention head (Figure 6 in Appendix), although we also test a few architectural variants in the number of decoder layers ($\mathsf{L}$) and number of attention heads per layer ($\mathsf{H}$) for comparisons (see Section 3 for details). The decoder layer contains masked self-attention so that for each position in the sequence the model can only attend to the current and previous positions. We choose to omit multi-layer feed-forward networks (FFNs) and layer normalization in the original Transformer model to examine the role of self-attention directly without interference from complex internal transformations introduced by FFNs and layer norms. The decoder layer is followed by an unembedding layer to project the decoder outputs to two logits (representing match and nonmatch) for each position.
@@ 3 Results
Figure 3: the model learns to attend target locations over training epochs. Here we show attention maps of a 1-layer 1-head Transformer model trained on the 3-back task as an example. See Appendix for attention maps in the 1-back and 2-back tasks.
@@
Figure 4: (a)-(c): the relationship between test accuracy at position $i$ and the attention score at position $i-N$ for the 1-layer 1-head Transformer model. Different colors represent different training epochs each dot belongs to. (d)-(f): the relationship between test accuracy at position $i$ and the attention score at position $i-N$ for the 1-layer 1-head Transformer model, but here different colors indicate different positions in the task sequence.
@@
To further investigate the relationship between attention scores and test accuracy, we plot accuracy at position $i$ against the attention score at the position $i-N$ over training epochs ($i\in\llbracket 1,24\rrbracket$, $N\in\{1,2,3\}$). The accuracy at position $i$ is defined as the percentage of the model making a correct prediction at position $i$. Over training epochs, we find that the attention score at position $i-N$ increases along with the accuracy at position $i$ (Figure 4a-c), and this is particularly observable for a large $N$ ($N\geq 2$). We reason that in order to produce an accurate prediction at position $i$, the Transformer model needs to learn to put most attention on the $i-N$ position and reduce dispersion of attention to other positions. To better visualize dispersion of attention scores across positions, we use the same data in Figure 4a-c but assign colors to the dots according to which position each dot belongs to (Figure 4d-f). This reveals a clear pattern that attention scores get dispersed at later locations, suggesting that more interference is caused when there are more preceding positions.
@@
For the 1-layer 1-head model, we find that $H_{N}$ increases as N increases, leading to the decrease in test accuracy (Figure 5). We infer that as N increases, it gets harder for the model to learn to attend to the N-back letter and the model is less confident about which letter is important, leading to higher entropy and lower accuracy. The fact that large values of $N$ require more structured attention weights (small entropy) to generalize in the N-back task is consistent with previous studies on representational learning theory [liu2022towards].
4 Discussion
@@
Although it is still unclear how selective attention in the human brain works at the algorithmic level, we can possibly draw inspirations from how the brain trades off between the amount vs. precision of the information being processed and design better model architectures with enhanced working memory capacity, which could in-turn lead to more powerful model capabilities in reasoning and problem solving [jaeggiImprovingFluidIntelligence2008, halfordSeparatingCognitiveCapacity2007].
Note that the current study focuses on a very simplified version of the Transformer model, so it is not straightforward to draw direct comparisons with pre-trained LLMs such as those evaluated by Gong et al. [9]. It is thus important for future research to investigate how the complexity of the model architecture and the number of learnable parameters in the model would influence task performance. In addition, varying the amount of training data and the specific hyperparameters used during training would also be crucial for understanding model behaviors in finer detail.